OpenStudy (anonymous):

(Line Integral) integrate F over the given curve: f(x,y)=x^3/y C: y=x^2/2 0<=x<=2

OpenStudy (anonymous):

my problem is coming up with r(t)

OpenStudy (anonymous):

i dont really understand what you want. the equation of the curve is y=0.5*x^2. the intergration needs to take place between x = 0 and 2. what does the function have to do with anything?

OpenStudy (anonymous):

it is a line integral over a given curve related to vector fields

OpenStudy (anonymous):

Calculus 3

OpenStudy (anonymous):

its 5 years since ive done calculus, and i dont integrate on a daily basis anymore, so i cant remember everything anymore..hope someone else can help

OpenStudy (anonymous):

r(t)=t^2/2 with 0<=t<=2

OpenStudy (anonymous):

x=t, y=t^2/2

OpenStudy (anonymous):

If you're desperate about using t anyway.

OpenStudy (anonymous):

THomas I thought so too, but I can't get the right answer. I think it has to do with Cos t and Sin t

OpenStudy (anonymous):

I don't think it's sine or cosine, because the curve isn't a circle or an ellipse of some kind.

OpenStudy (anonymous):

can you work out the problem with x=t y=.5t^2 and see if we get smae answer. The anser is 10(5^1/2)-2/3

OpenStudy (anonymous):

i get something totally different

OpenStudy (anonymous):

You get: \[\int\limits_{0}^{2}2tdt\] right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

2root2

OpenStudy (anonymous):

you have to take vector and multiply |v|

OpenStudy (anonymous):

f(x,y)=int a to b (g(t,h(t))|v(t)|dt

OpenStudy (anonymous):

I don't know how this works anymore I'm afraid. I think you're right that the integrand isn't just 2t, but I'm not sure what to add. What do you mean by |v(t)|?

OpenStudy (anonymous):

|r '(t)|?

OpenStudy (anonymous):

if v=ti+tj+0k v=sqrt 1^2+1^2= sqrt 2=|v|

OpenStudy (anonymous):

\[|v|=\sqrt{2t^{2}}=\sqrt{2}t\] right? but what's v in this context?

OpenStudy (anonymous):

v is the length of the curve from my understanding, in the formula |v(t)| is a constant that we take out of the integral and multiply at the end. It is derived from the vector. I got the same as you, but I cannot figure out how they arrived at 10root 5 something!

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=line+integral&lk=1&a=ClashPrefs_*MathWorld.LineIntegral- This made things clearer for me.

OpenStudy (anonymous):

Wait, we don't have a vector field, we have a scalar field.

OpenStudy (anonymous):

Yes scalar, vector field is the next topic, but I have to get a handle on this first :)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=path+integral This applies.

OpenStudy (anonymous):

yes that is the formula

OpenStudy (anonymous):

So we need to compute |r '(t)|.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

r(t)=(t,t^2/2), so what's r '(t)?

OpenStudy (anonymous):

1^2 1/2^2?sqroot?

OpenStudy (anonymous):

or is it 1^2 1^2=sqrt2?

OpenStudy (anonymous):

wait is that r prime t?

OpenStudy (anonymous):

yes, the derivative.

OpenStudy (anonymous):

my book gives me V(t)

OpenStudy (anonymous):

not prime

OpenStudy (anonymous):

what's prime?

OpenStudy (anonymous):

derivative no?

OpenStudy (anonymous):

Right, we need to compute the derivative of r(t), which is a vector. Just derive both components to t.

OpenStudy (anonymous):

oh crap Velocity is the derivative of R(t) isn't it.

OpenStudy (anonymous):

Oh, probably.

OpenStudy (anonymous):

and acceleration is double prime of r(t) or prime of V

OpenStudy (anonymous):

Even if we accept V(t) is the derivative, I still do not get the right answer though.

OpenStudy (anonymous):

derivative would be 1^2 +1^2 all under sqroot= sqrt 2 yes?

OpenStudy (anonymous):

v(t)=(1,t)

OpenStudy (anonymous):

r(t)=(t,t^2/2), so for v(t) you derive both components to t.

OpenStudy (anonymous):

er, right sqrt 1+t how does that help?

OpenStudy (anonymous):

\[\sqrt{1+t^{2}}\]

OpenStudy (anonymous):

yep again i erred, so we tack it onto the end of t^3/t^2?

OpenStudy (anonymous):

yes, so we get \[\int\limits_{0}^{2}2t\sqrt{1+t^{2}}\]

OpenStudy (anonymous):

im with you so far.

OpenStudy (anonymous):

integration by parts?

OpenStudy (anonymous):

no u substitution

OpenStudy (anonymous):

that's right.

OpenStudy (anonymous):

\[\int\limits_{0}^{\sqrt{2}}\sqrt{1+u}du\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

no 1/2sqrtu

OpenStudy (anonymous):

from 1 to 5

OpenStudy (anonymous):

let me work this out i think your on right track

OpenStudy (anonymous):

\[2\int\limits_{0}^{2}t ^{3}/t ^{2}\sqrt{1+t ^{2}}\] yes?

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

My substitution is wrong, I'm not sure why.

OpenStudy (anonymous):

ll =1 ul=5 u sub =1+t^2 dr us ub =2t

OpenStudy (anonymous):

ll=0+1=1 ul=2^2+1=5

OpenStudy (anonymous):

I just worked out the whole problem Its correct!!!! THANK YOU!!!!

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

I've been working on this crap for 2 days, taking a break good job mate!

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/Classes/CalcIII/CalcIII.aspx Here is a good Calc 3 supplement.